OpenStudy (anonymous):

Which is the equation of the given line? A. x = –2 B. x = 2 C. y = –2 D. y = –2x

OpenStudy (anonymous):

@green_1

OpenStudy (anonymous):

james1769

OpenStudy (anonymous):

@james1769

OpenStudy (green_1):

i think it is y=-2

OpenStudy (anonymous):

Ok thanks

OpenStudy (green_1):

any other questions

OpenStudy (anonymous):

yeah 1 more

OpenStudy (anonymous):

cant copy it

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

Will give medal

OpenStudy (trojanpoem):

y = -2

OpenStudy (trojanpoem):

Simply for any value given for x , y remains constant ( -2 )

OpenStudy (green_1):

i think the answer is "C" because the y which is 2 in the equation hits 2 on the graph and the slope -3 maches the graph be cause the points on the graph goes down three and across 1

OpenStudy (trojanpoem):

@green_1 , Isn't the slope zero ? \[\frac{ y_{2} - y_{1} }{ x_{2} - x_{1} } = \frac{ -2 - (-2) }{ x_{2} - x_{1} } = 0\]

OpenStudy (green_1):

the equations are already in slope intercept form. all you have to do is see where the line on the graph hits the y axis, which is 2 and to find the slope all you have to do is use rise over run when you are following the points showed and the points in between them.

OpenStudy (green_1):

i think you would use that equation on finding the equation for ordered pairs.

OpenStudy (trojanpoem):

"to find the slope all you have to do is use rise over run when you are following the points showed and the points in between them." The slope is zero as the line is horizontal.

OpenStudy (trojanpoem):

@green_1 , You're right, he could use the equation to find the it. y = mx + b we found that slope = 0 , y intersection is -2 b = -2 m = 0 so y = -2